### v2

@@ Abstract
We consider the problem of preprocessing an $n\times n$ matrix $\mathbf{M}$, and supporting queries that, for any vector $v$, returns the matrix-vector product $\mathbf{M}v$. This problem has been extensively studied in both theory and practice: on one side, practitioners have developed algorithms that are highly efficient in practice, whereas on the other side, theoreticians have proven that the problem cannot be solved faster than naive multiplication in the worst-case. This lower bound holds even in the average-case, implying that existing average-case analyses cannot explain this gap between theory and practice. Therefore, we study the problem for structured matrices. We show that for $n\times n$ matrices of VC-dimension $d$, the matrix-vector multiplication problem can be solved with $\widetilde{O}(n^{2})$ preprocessing and $\widetilde{O}(n^{2-1/d})$ query time. Given the low constant VC-dimensions observed in most real-world data, our results posit an explanation for why the problem can be solved so much faster in practice. Moreover, our bounds hold even if the matrix does not have a low VC-dimension, but is obtained by (possibly adversarially) corrupting at most a subquadratic number of entries of any unknown low VC-dimension matrix.
In turn, our results also yield the first non-trivial upper bounds for many applications. In previous works, the online matrix-vector (OMv) hypothesis (conjecturing that quadratic time is needed per query, even over the boolean semi-ring) was used to prove many conditional lower bounds, showing that it is impossible to compute and maintain high-accuracy estimates for shortest paths, Laplacian solvers, effective resistance, and triangle detection in graphs subject to node insertions and deletions in subquadratic time. Yet, via a reduction to our matrix-vector-multiplication result, we show we can maintain the aforementioned problems efficiently if the input is structured, providing the first subquadratic upper bounds in the high-accuracy regime.
keywords:
Online Matrix-Vector Multiplication, Boolean Matrix Multiplication, VC dimension, Pollard Pseudodimension, Parameterized Complexity, Dynamic Laplacian Solver
footnotetext: Work done while at Princeton University.
@@ 1 Introduction
Computing sequential matrix-vector products is a workhorse of many iterative algorithms in machine learning. In optimization (Renegar, 1996; Vempala et al., 2020; Lin et al., 2023a; Chaudhari et al., 2024), computational geometry (Har-Peled, 2009; Welzl, 1992; Fisikopoulos and Penaranda, 2016), online algorithms (Lin et al., 2024; Anand and Qu, 2024; Murray et al., 2021), and dynamic algorithms (Liu, 2024; Jiang et al., 2023; Anand et al., 2024b; Jin and Xu, 2022; Lin et al., 2023b), sequential matrix-vector products are essential subroutines underlying performant algorithms. The current learning revolution is powered by hardware specifically designed to perform such products, as both neural network evaluation and back-propagation require repeated matrix-vector products (Sørensen, 2012; Rumelhart et al., 1986). Hence, matrix-vector products are arguably one of the most important and fundamental subroutines, with any complexity improvement having wide-ranging implications, and is thus a prevalent research topic in both theory and practice.
@@
The problem can be modeled via the following data structure task: construct a data structure that preprocesses a given $n\times n$ matrix $\mathbf{M}$. After preprocessing, we want to multiply vectors with $\mathbf{M}$ faster than naive matrix-vector multiplication. Depending on the application, this matrix $\mathbf{M}$ is fixed or the data structure may need to handle changes to $\mathbf{M}$. For instance, when calculating probabilities of a random walk (e.g., in the PageRank algorithm (Page et al., 1999)), performing the power-method, and evaluating a neural network, the matrix is fixed. On the other hand, in convex optimization problems, such as in solving a linear or semi-definite program, applying Newton-Raphson’s method, and training a neural network, the matrix is either the inverse of a Hessian, or is given by the network’s weights, which changes from one iteration to the next.
@@
Beyond speedups due to improving hardware accelerators, practitioners have made tremendous progress in accelerating the computation of matrix-vector products through heuristics that run in $\mathrm{nnz}(\mathbf{M})$ (the number of non-zero entries of $\mathbf{M}$) worst-case time, but which are much faster in practice. This has led to speedups in training graph neural networks (Alves et al., 2024) and other learning algorithms (Floros et al., 2024).
@@
While practitioners have had success in accelerating the computation of matrix vector products, from a theory perspective there are substantial lower bounds. Gronlund and Larsen (2015) showed that any $\mathrm{poly}(n)$-space data structure for arbitrary matrix-vector multiplication over sufficiently large fields (such as $\mathbb{R}$) require $\Omega(n^{2}/\log n)$ time. In particular, any “generic” algorithm that does not make assumptions about the specific field needs $\Omega(n^{2}/\log n)$ time. This quadratic lower bound was also proven for arithmetic circuits (Frandsen et al., 2001). While these are worst-case lower bounds, they also hold for the average case (Henzinger et al., 2022). The only non-trivial upper bounds beating quadratic time are over the Boolean semi-ring (i.e., $\{0,1\}$ with $x\oplus y=\min(1,x+y)$). Here, the famous mailman algorithm (Liberty and Zucker, 2006) achieves $O(n^{2}/\log n)$ runtime. In a line of work that reduced the problem to finding neighborhoods in a graph and smaller algebraic products, this query complexity was improved to $O(n^{2}/\log^{2}n)$ (Williams, 2007) and $O(n^{2-o(1)})$ (Larsen and Williams, 2017; Chakraborty et al., 2018; Abboud et al., 2024). However, none of these algorithms are truly subquadratic, i.e., $O(n^{2-\epsilon})$ for some constant $\epsilon>0$.
@@
Moreover, while the $\Omega(n^{2})$ lower bounds only hold for fields, even over the Boolean semi-ring it is conjectured that no truly subquadratic time algorithm exists (Henzinger et al., 2015), referred to as the OMv (online matrix-vector multiplication) conjecture. This has led to a line of works that use the conditional hardness of OMv to prove tight time lower bounds for a host of dynamic algorithms, such as dynamic matrix inversion (Brand et al., 2019a, b), dynamic subgraph connectivity (Henzinger and Neumann, 2016), dynamic regression (Jiang et al., 2023), dynamic range mode (Jin and Xu, 2022), Erickson’s problem and Langerman’s problem (Jin and Xu, 2022), as well as high-dimensional generalizations of Klee’s measure problem (Yildiz and Suri, 2012).
@@
In summary, there is a substantial difference between the perspectives of this problem in theory and practice: there are heuristics that are fast in practice, but from a theoretical perspective the problem is hard, and an entire subarea of fine-grained-complexity has been built on this hardness. Moreover, average-case analysis cannot explain this difference between theory and practice, as even the average-case is provably hard. In this work, we resolve this conflict. Since the average-case (i.e., random non-structured input) is hard, the observed efficiency of practical algorithms must stem from some inherent structure in the real-world input data. Thus, we must study the complexity for structured inputs.
@@
A popular measure for structural complexity is the Vapnik-Chervonenkis (VC) dimension, which finds many applications in machine learning (Shalev-Shwartz and Ben-David, 2014; Bartlett et al., 2019), structural graph theory (Nguyen et al., 2024; Alon et al., 2016; Karczmarz and Zheng, 2024), computational geometry (Har-Peled, 2009; B. Chazelle, 1989; Fisikopoulos and Penaranda, 2016) (see the preliminaries in Section 2 for a definition of the VC-dimension). Many structured objects studied by theoreticians have inherently low VC dimension. For instance, the hypothesis class of intervals on $\mathbb{R}$, star and interval graphs, and rank-one projection matrices, have VC-dimension $2$, planar graphs have VC dimension 4, and any $H$-minor free graph has VC dimension at most $|V(H)|-1$. Beyond being an established complexity measure in theory, it has also been empirically verified that real-world data has low VC dimension (Coudert et al., 2024). This observation can be explained via the following structural characterization of matrices with low VC dimension:
[…]
@@
Presumably, whatever unknown structure is common in real-world matrices should not be lost when deleting rows/columns. We note that section 1 does not preclude other matrices, that do not satisfy this hereditary property, from having constant VC dimension, e.g., adjacency matrices of minor-free graphs.
@@ 1.1 Our Results
While theoretical objects such as grid-graphs, intersection graphs, and kernel-matrices have a low VC-dimension, they are not perfect representations of real-world data. For instance, while city street networks in the US are predominantly grid-graphs, there are always exceptions in form of bridges or a few diagonal streets. Exceptions can also occur due to errors in measurements or other corruptions. To capture such almost low VC-dimension objects, we define the “corrupted VC-dimension.”
[…]
We say that a set system $\mathcal{F}$ on a ground set of size $n$ has “corrupted VC-dimension $d$” if there is another set system $\mathcal{F}^{\prime}$ of VC-dimension $\leq d$ such that $\mathcal{F}$ can be obtained from $\mathcal{F}^{\prime}$ by adding and/or removing (i.e., corrupting) at most $O(n^{1-1/d})$ elements to/from each set.
For matrices $\mathbf{M}\in\{0,1\}^{m\times n}$, the corrupted VC-dimension $d$ is equivalent to the existence of matrices $\mathbf{L}\in\{0,1\}^{m\times n}$ and $\mathbf{S}\in\{-1,0,1\}^{m\times n}$ with $\mathbf{M}=\mathbf{L}+\mathbf{S}$ where $\mathbf{L}$ has VC-dimension $d$ and the corruption matrix $\mathbf{S}$ has at most $O(n^{1-1/d})$ non-zero entries per row, i.e., flipping at most $O(n^{1-1/d})$ bits per row of $\mathbf{L}$.
Note that merely the existence of such an $\mathcal{F}^{\prime}$ (i.e., $\mathbf{L}$) of VC-dimension $d$ suffices for $\mathcal{F}$ (i.e., $\mathbf{M}$) to have corrupted VC-dimension $d$. We do not need to know $\mathcal{F}^{\prime}$ (i.e., $\mathbf{L}$) or which sets have been corrupted (i.e., $\mathbf{S}$). Furthermore, set families of corrupted VC-dimension $d$ are strict generalizations of set families of VC-dimension exactly $d$. Therefore, while we state our main results for matrices with corrupted VC-dimension $d$, they also hold for general matrices with VC-dimension $d$.
We now state the theoretical guarantees of our accelerated matrix-vector multiplication algorithm in Theorem 1.1, where the runtimes are parameterized by the corrupted VC dimension measure.
@@
(Static Online Matrix-Vector Multiplication). If a matrix $\mathbf{M}\in\{0,1\}^{m\times n}$ has corrupted VC-dimension $d$, then after an $\widetilde{O}(mn)$-time preprocessing, there is a data structure $\mathcal{D}$ that can compute $\mathbf{M}v$ for any $v\in\mathbb{R}^{n}$ in $\widetilde{O}(mn^{1-1/d}+n)$ time, with high probability.
The problem of approximating the VC-dimension of any set system within a $(2-\epsilon)$ accuracy for any $\epsilon>0$ is known to be $\Sigma_{3}^{P}$-hard (Mossel and Umans, 2002); however, for our theorem, the algorithm does not need to know or compute the VC-dimension to achieve this runtime.
This data structure is based on Gasieniec and Lingas (2003) which proposed an algorithm for Boolean matrix multiplication. Their algorithm computes the matrix-matrix product via $n$ matrix-vector products, thus implicitly providing a matrix-vector multiplication data structure as well. Their complexity depends on the weight of the minimum spanning tree defined with respect to the Hamming-distance between the rows of $\mathbf{M}$. While $O(n^{2})$ in worst-case, they conjecture this should be unlikely in real-world data. This data structure was recently independently rediscovered in the graph neural networks community (Alves et al., 2024) where they experimentally verified that the algorithm is indeed efficient on real-world inputs. Our Theorem 1.1 now gives a theoretical explanation as to why these practical algorithm behave so much more efficient in practice than the worst-case and average-case lower bounds of $\Omega(mn)$.
We obtain Theorem 1.1 via techniques from computational geometry. A line of works (Welzl, 1988; B. Chazelle, 1989; Matoušek, 1991; Har-Peled, 2009) consider the following geometric intersection data structure problem: given a set of points in $\mathbb{R}^{d}$, preprocess them. Then, given a convex polytope for each query, return whether any of the points intersect the polytope.
Matrix vector products can be interpreted as $O(m)$ intersection problems if we only care about Boolean outputs, because in the Boolean case $(\mathbf{M}v)_{i}\neq 0$ if and only if there is an intersection between the position of the non-zero elements in $v$ and the non-zero elements in the $i$-th row of $\mathbf{M}$. We show that the computational geometry techniques work even for non-boolean outputs, and we also extend them from bounded VC-dimension to our notion of corrupted VC-dimension. We provide details of this proof in Section 3.2.
At last, we remark that prior work for the geometric intersection problem had exponential preprocessing time (Welzl, 1992) or unspecified polynomial time (Matoušek, 1991).
In addition to our novel complexity bound, we also provide alternative algorithms to those presented in Gasieniec and Lingas (2003); Alves et al. (2024). We provide more details below.
Dynamic Setting
We extend Theorem 1.1 by allowing $\mathbf{M}$ to undergo row and column updates. The proof is given in Section A.2.
@@
(Dynamic Online Matrix-Vector Multiplication). Given a matrix $\mathbf{M}\in\{0,1\}^{m\times n}$, there is a data structure $\mathcal{D}$ with $\widetilde{O}(mn)$ preprocessing time that supports row and column updates (insertions/deletions) to $\mathbf{M}$ in $\widetilde{O}(n)$ and $\widetilde{O}(m)$ time respectively. Upon querying the data structure $\mathcal{D}$ with a vector $v\in\mathbb{R}^{n}$, it outputs $\mathbf{M}v$ in $\widetilde{O}(mn^{1-1/d^{*}}+n)$ time, with high probability, where $d^{*}$ is the largest corrupted VC-dimension of $\mathbf{M}$ throughout the history of its updates.
In the special case where the updates to $\mathbf{M}$ keep its corrupted VC-dimension $d$ invariant, each matrix-vector product is computed with high probability in $\widetilde{O}(mn^{1-1/d}+n)$ time.
Extensions to Transposed Matrices.
Assume we have a small VC-dimension bound on $\mathbf{M}^{\top}$, then the VC-dimension of $\mathbf{M}$ could be much larger than that of $\mathbf{M}^{\top}$. If $d$ bounds the VC-dimension of $\mathbf{M}^{\top}$, then the best bound on the VC-dimension of $\mathbf{M}$ is $2^{d+1}$ (Assouad, 1983), so Theorem 1.1 would have only $\widetilde{O}(mn^{1-1/2^{d}})$ query time. To alleviate this substantial slow-down in Theorems 1.1 and 1.2, we extend the algorithms so their complexity is also bounded by the VC-dimension of $\mathbf{M}^{\top}$.
When $\mathbf{M}^{\top}\in\{0,1\}^{n\times m}$ has corrupted VC-dimension $d^{\prime}$, we prove analogous versions of Theorems 1.1 and 1.2 that allow computing $\mathbf{M}v$ with high probability in $O(nm^{1-1/d^{\prime}}+m)$ time after a $\widetilde{O}(mn)$-time preprocessing. Thus, if $\mathbf{M}\in\{0,1\}^{m\times n}$ has corrupted VC-dimension $d$ and $\mathbf{M}^{\top}\in\{0,1\}^{n\times m}$ has corrupted VC-dimension $d^{\prime}$, then after an $\widetilde{O}(mn)$-time preprocessing, we can compute $\mathbf{M}v$ in time $\min\{O(mn^{1-1/d}+n,nm^{1-1/d^{\prime}}+m)\}$. For this, we do not need to know $d$ or $d^{\prime}$, as the algorithm automatically achieves the minimum runtime.
Extension to Non-Boolean Matrices.
The Pollard pseudodimension is a popular extension of the VC-dimension to non-binary thresholds, which also finds many applications in theoretical machine learning (Shalev-Shwartz and Ben-David, 2014; Bartlett et al., 2019) and characterizations in structural graph theory (Karczmarz and Zheng, 2024) (see the preliminaries in Section 2 for a definition of the Pollard pseudodimension). For a real-valued matrix $\mathbf{M}\in\{0,1\}^{m\times n}$, if $d$ bounds the Pollard pseudodimension of $\mathbf{M}$ and if $A$ bounds the number of thresholds of $\mathbf{M}$, then by lemma A.13, $\mathbf{M}v$ can be computed in time $\widetilde{O}(Amn^{1-1/d}+n)$. For $A\leq\widetilde{O}(1)$, this provides a subquadratic runtime.

### v3

@@ Abstract
We consider the problem of preprocessing an $n\times n$ matrix $\mathbf{M}$, and supporting queries that, for any vector $v$, returns the matrix-vector product $\mathbf{M}v$. This problem has been extensively studied in both theory and practice: on one side, practitioners have developed algorithms that are highly efficient in practice, whereas on the other side, theoreticians have proven that the problem cannot be solved faster than naive multiplication in the worst-case. This lower bound holds even in the average-case, implying that existing average-case analyses cannot explain this gap between theory and practice. Hence, we study the problem for structured matrices. We show that for $n\times n$ Boolean matrices of VC-dimension $d$, the matrix-vector multiplication problem can be solved with $\smash{\widetilde{O}(n^{2})}$ preprocessing and $\smash{\widetilde{O}(n^{2-1/d})}$ query time. Given the low constant VC-dimensions observed in most real-world data, our results posit an explanation for why the problem can be solved so much faster in practice. Furthermore, we show how to extend this result to the non-Boolean setting with the Pollard pseudodimension.
Our results yield the first non-trivial upper bounds for many applications. In previous works, the online matrix-vector (OMv) hypothesis (conjecturing that quadratic time is needed per query, even over the boolean semi-ring) was used to prove many conditional lower bounds, showing that it is impossible to compute and maintain high-accuracy estimates for effective resistance, Laplacian solvers, shortest paths, and triangle detection in graphs subject to node insertions and deletions in subquadratic time. Yet, via a reduction to our matrix-vector-multiplication result, we show we can maintain these problems efficiently if the input is structured, providing the first subquadratic upper bounds in the high-accuracy regime.
@@ 1 Introduction
Computing sequential matrix-vector products is a fundamental subroutine of many iterative algorithms in machine learning. In optimization (Renegar, 1996; Vempala et al., 2020; Lin et al., 2023b; Chaudhari et al., 2024), computational geometry (Har-Peled, 2009; Welzl,Emo, 1992; Fisikopoulos and Penaranda, 2016), online algorithms (Lin et al., 2024; Anand and Qu, 2024; Murray et al., 2021), and dynamic algorithms (Liu, 2024; Jiang et al., 2023; Anand et al., 2024b; Jin and Xu, 2022), sequential matrix-vector products are essential subroutines underlying performant algorithms. The current learning revolution is powered by hardware specifically designed to perform such products, as both neural network evaluation and back-propagation require repeated matrix-vector products (Sørensen, 2012; Rumelhart et al., 1986). Hence, matrix-vector products are arguably one of the most important and fundamental subroutines, with any complexity improvement having wide-ranging implications, and is thus a prevalent research topic in both theory and practice.
@@
The problem can be modeled via the following data structure task: construct a data structure that preprocesses a given $n\times n$ matrix $\mathbf{M}$. After preprocessing, we want to multiply vectors with $\mathbf{M}$ faster than naive matrix-vector multiplication. Depending on the application, this matrix $\mathbf{M}$ is fixed or the data structure may need to handle changes to $\mathbf{M}$. For instance, when calculating probabilities of a random walk (e.g., in the PageRank algorithm (Page et al., 1999)), performing the power-method, and evaluating a neural network, the matrix is fixed. Conversely, in convex optimization, such as when solving a linear or semi-definite program, applying Newton-Raphson’s method, and training a neural network, the matrix is the inverse of a Hessian or is given by the network’s weights, which changes from one iteration to the next (Arriaga and Vempala, 2006; Anand and Liaw, 2025).
@@
Beyond speedups due to improved hardware, practitioners have made tremendous progress in accelerating the computation of matrix-vector products through heuristics that run in $\mathrm{nnz}(\mathbf{M})$ (the number of non-zero entries of $\mathbf{M}$) worst-case time, but which are much faster in practice. This has led to speedups in training GNNs (Alves et al., 2024) and other algorithms (Floros et al., 2024).
@@
Conversely, from a theory perspective there are substantial lower bounds. Gronlund and Larsen (2015) showed that any $\mathrm{poly}(n)$-space data structure for matrix-vector multiplication over sufficiently large fields (for instance, $\mathbb{R}$) require $\Omega(n^{2}/\log n)$ time. This quadratic lower bound was also proven for arithmetic circuits (Frandsen et al., 2001). While these are worst-case lower bounds, they also hold for the average case (Henzinger et al., 2022). The only non-trivial upper bounds beating quadratic time are over the Boolean semi-ring (i.e., $\{0,1\}$ with $x\oplus y=\min(1,x+y)$). In a line of works that reduced the problem to finding neighborhoods in a graph and smaller algebraic products, this query complexity was improved to $O(n^{2}/\log^{2}n)$ (Liberty and Zucker, 2006; Williams, 2007) and $O(n^{2-o(1)})$ (Larsen and Williams, 2017; Chakraborty et al., 2018; Abboud et al., 2024). However, none of these algorithms are truly subquadratic, i.e., $O(n^{2-\epsilon})$ for some constant $\epsilon>0$.
@@
Moreover, while the $\Omega(n^{2})$ lower bounds only hold over finite fields, even over the Boolean semi-ring it is conjectured that no truly subquadratic time algorithm exists (Henzinger et al., 2015), and this is commonly referred to as the OMv (online matrix-vector multiplication) conjecture. In fact, this has led to a recent line of works that use the conditional hardness of OMv to prove tight time lower bounds for many dynamic algorithms, such as dynamic matrix inversion (Brand et al., 2019a, b), dynamic subgraph connectivity (Henzinger and Neumann, 2016), dynamic regression (Jiang et al., 2023), dynamic range, Langerman’s problem (Jin and Xu, 2022), and the generalized Klee’s high-dimensional measure problem (Yildiz and Suri, 2012).
@@
In summary, there is a notable difference between the perspectives of this problem in theory and practice: there are heuristics that are fast in practice, but from a theoretical perspective the problem is hard, and an entire subarea of fine-grained-complexity has been built on this hardness assumption (Hu and Polak, 2024). Further, average-case analyses cannot explain this gap, as even the average-case is provably hard. In this work, we resolve this conflict: since the average-case (i.e., random non-structured inputs) is hard, the observed efficiency of practical algorithms must stem from some inherent structure in real-world data. Thus, we study the complexity for structured inputs.
@@
A popular measure for structural complexity is the Vapnik-Chervonenkis (VC) dimension, which finds many applications in machine learning (Shalev-Shwartz and Ben-David, 2014; Bartlett et al., 2019), structural graph theory (Nguyen et al., 2024; Alon et al., 2016; Karczmarz and Zheng, 2024), and computational geometry (Har-Peled, 2009; B. Chazelle, 1989; Fisikopoulos and Penaranda, 2016) (see the preliminaries in Section˜2 for a definition of the VC-dimension). For Boolean matrices, the VC-dimension has the following, more intuitive, description: it is the size of the largest subset $S\subseteq[n]$ of columns such that the rows of the sub-matrix whose columns are restricted to those in $S$ contain every possible string in $\{0,1\}^{|S|}$. Many structured objects studied by theoreticians have inherently low VC dimension. For instance, star and interval graphs, and rank-one projection matrices, have VC-dimension $2$, planar graphs have VC dimension 4, and any $H$-minor free graph has VC dimension at most |V(H)|1|V(H)|-1. Beyond being an established complexity measure in theory, it has also been empirically observed that real-world data has low VC dimension (Coudert et al., 2024). We explain this observation via the following structural characterization of matrices with low VC dimension:
Theorem 1.1 (Proof in Appendix˜D).
@@
Presumably, any unknown structure common to real-world matrices should not be lost when deleting rows/columns. Theorem˜1.1 does not exclude other matrices (which do not satisfy this hereditary property) from having constant VC-dimension, e.g., adjacency matrices of minor-free graphs.
@@ 1.1 Our Results
While theoretical objects such as grid-graphs, intersection graphs, and kernel-matrices have a low VC-dimension, they are not perfect representations of real-world data. For instance, while city street networks are predominantly grid-graphs, there always exists exceptions in the form of bridges or few diagonal streets. Exceptions can also occur due to errors in measurements or other corruptions. To capture such almost low VC-dimension objects, we define the “corrupted VC-dimension.”
Definition 1.2 (Corrupted VC-dimension d).
The “corrupted VC-dimension” of a set system $\mathcal{F}$ with $m$ sets is the smallest $d$ such that there is another set system $\mathcal{F}^{\prime}$ of VC-dimension $\leq d$ and $\mathcal{F}$ can be obtained from $\mathcal{F}^{\prime}$ by adding and/or removing each element to/from at most $O(m^{1-1/d})$ sets.
For matrices $\mathbf{M}\in\{0,1\}^{m\times n}$, a corrupted VC-dimension $d$ implies the existence of a matrix $\mathbf{L}\in\{0,1\}^{m\times n}$ of VC-dimension $\leq d$ and a ‘corruption’ matrix $\mathbf{S}\in\{-1,0,1\}^{m\times n}$ with at most $O(m^{1-1/d})$ non-zero entries per column, such that $\mathbf{M}=\mathbf{L}+\mathbf{S}$. Note that the mere existence of such an $\mathcal{F}^{\prime}$ (i.e., $\mathbf{L}$) of VC-dimension $\leq d$ suffices for $\mathcal{F}$ (i.e., $\mathbf{M}$) to have corrupted VC-dimension $d$. We do not need to know $\mathcal{F}^{\prime}$ or which sets have been corrupted. Also, the VC-dimension of $\mathcal{F}$ upper bounds the corrupted VC-dimension of $\mathcal{F}$. Hence, while we state our results for matrices with corrupted VC-dimension $d$, the proven upper bounds also hold for matrices with VC-dimension $d$.
We now state the theoretical guarantees of our accelerated matrix-vector multiplication algorithm in Theorem˜1.3, where the runtimes are parameterized by the corrupted VC dimension measure.
@@
(Static Online Matrix-Vector Multiplication). If a matrix $\mathbf{M}\in\{0,1\}^{m\times n}$ has corrupted VC-dimension $d$, then after an $\widetilde{O}(mn)$-time preprocessing, there is a data structure $\mathcal{D}$ that can compute $\mathbf{M}v$ for any $v\in\mathbb{R}^{n}$ in $\widetilde{O}(nm^{1-1/d}+m)$ time, with high probability.
The problem of approximating the VC-dimension of any set system within a $(2-\epsilon)$ accuracy for any $\epsilon>0$ is known to be $\Sigma_{3}^{P}$-hard (Mossel and Umans, 2002); however, for our theorem, the algorithm does not need to know or compute or even approximate the VC-dimension to achieve this runtime.
Our data structure is based on Björklund and Lingas (2001) which proposed an algorithm for Boolean matrix multiplication. Their algorithm computes the matrix-matrix product via $n$ matrix-vector products, thus implicitly providing a matrix-vector multiplication data structure as well. Their complexity depends on the weight of the minimum spanning tree (MST) defined with respect to the Hamming-distance between the rows of $\mathbf{M}$. However, the weight of the MST is $O(n^{2})$ in the worst-case.
This algorithm was recently independently rediscovered in the graph neural networks community (Alves et al., 2024) which experimentally verified its efficiency on real-world inputs. Our Theorem˜1.3 now gives a theoretical explanation as to why these practical algorithms are much more efficient in practice than the worst-case and average-case lower bounds of $\Omega(mn)$.
We obtain Theorem˜1.3 via techniques from computational geometry. A line of works (Welzl, 1988; B. Chazelle, 1989; Matoušek1991; Har-Peled, 2009) consider the following geometric intersection data structure problem: given a set of points in $\mathbb{R}^{d}$, preprocess them. Then, given a convex polytope for each query, return whether any of the points intersect the polytope. However, prior work had exponential preprocessing time (Welzl,Emo, 1992) or unspecified polynomial time (Matoušek1991). Matrix vector products can be interpreted as $O(m)$ intersection problems if we only care about Boolean outputs, because in the Boolean case $(\mathbf{M}v)_{i}\neq 0$ if and only if there is an intersection between the position of the non-zero elements in $v$ and the non-zero elements in the $i$-th row of $\mathbf{M}$. We extend these computational geometry techniques to non-Boolean outputs, and to our more general notion of corrupted VC-dimension. We provide details of this proof in Section˜3.2.
In addition to our novel complexity bound, we also provide alternative algorithms to those presented in Björklund and Lingas (2001); Alves et al. (2024). We provide more details below.
Dynamic Setting. We extend Theorem˜1.3 by allowing $\mathbf{M}$ to undergo row and column updates. The proof is given in Section˜B.3.
@@
(Dynamic Online Matrix-Vector Multiplication). Given a matrix $\mathbf{M}\in\{0,1\}^{m\times n}$, there is a data structure $\mathcal{D}$ with $\widetilde{O}(mn)$ preprocessing time that supports row and column updates (insertions/deletions) to $\mathbf{M}$ in $\widetilde{O}(n)$ and $\widetilde{O}(m)$ time, respectively. Upon querying $\mathcal{D}$ with a vector $v\in\mathbb{R}^{n}$, it outputs $\mathbf{M}v$ in $\widetilde{O}(nm^{1-1/d^{*}}+m)$ time, with high probability, where $d^{*}$ is the largest corrupted VC-dimension of $\mathbf{M}$ throughout the history of its updates.
Transposed Matrices. If the VC-dimension of $\mathbf{M}^{\top}$ is $d$, then the VC-dimension of $\mathbf{M}$ is at most $2^{d+1}$ (Assouad, 1983), so Theorem˜1.3 runs in $\widetilde{O}(nm^{1-1/2^{d}}+m)$ time. To address this slow-down, we extend the algorithm to bound its complexity by the VC-dimension of $\mathbf{M}^{\top}$.
If $\mathbf{M}\in\{0,1\}^{m\times n}$ and $\mathbf{M}^{\top}\in\{0,1\}^{n\times m}$ have corrupted VC-dimension $d$ and $d^{\prime}$ (respectively), then after an $\widetilde{O}(mn)$-time preprocessing, we show we can compute $\mathbf{M}v$ in time $\widetilde{O}(\min\{nm^{1-1/d}+m,mn^{1-1/d^{\prime}}+n\})$ with high probability. We do not need to know $d$ or $d^{\prime}$, as the algorithm automatically achieves the minimum.
Non-Boolean Matrices. The Pollard pseudodimension extends the VC-dimension to non-binary thresholds. It finds many applications in theoretical machine learning (Shalev-Shwartz and Ben-David, 2014; Bartlett et al., 2019) and structural graph theory (Karczmarz and Zheng, 2024) (see Section˜B.4 for a definition of the Pollard pseudodimension). For $\mathbf{M}\in\mathbb{R}^{m\times n}$, if $d$ bounds the Pollard pseudodimension of $\mathbf{M}$ and if $T$ bounds the number of unique values per column of $\mathbf{M}$, then by Theorem˜B.9, $\mathbf{M}v$ can be computed in time $\widetilde{O}(Tnm^{1-1/d}+m)$. For $T\leq\widetilde{O}(1)$, this is subquadratic.